# Okapi: Efficiently Safeguarding Speculative Data Accesses in Sandboxed Environments

### v2

@@
Abstract
This paper introduces Okapi, a new hardware/software cross-layer architecture designed to mitigate Transient Execution Side Channel (TES) attacks, including Spectre variants, in modern computing systems. Okapi enforces sandboxing for speculative execution, providing a hardware basis that can replace expensive speculation barriers in software.
@@ Abstract
Okapi introduces the OkapiReset instruction for additional software-level security support. This instruction allows for fine-grained sandboxing with custom program sizes smaller than a thread, resulting in 1.68 % performance overhead in our WebAssembly runtime experiment.
@@
1 Introduction
@@
Transient execution side channel (TES) attacks, including numerous variants of Spectre [28], persist as serious threats in modern computing systems. These vulnerabilities enable attackers to exploit side effects of speculative execution by tricking the processor into revealing information that is not accessed by the sequential semantics of the program. Since TES attacks exploit high-end features of modern microarchitectures such as speculative and out-of-order execution, eliminating them without introducing a significant performance penalty continues to be a major challenge. The most dangerous Spectre attacks involve so-called universal read gadgets that are capable of revealing the contents of any address in the entire memory space.
@@
A widely used software mechanism to restrict a program’s accessible resources is sandboxing [44]. Untrusted code is often recommended to be executed within a sandbox [21] in order to prevent it from stealing data or otherwise compromising the system [36]. As long as no additional measures are in place, Spectre-style attacks pose a severe threat to such software-level isolation techniques [40]. Attackers can either perform a breakout attack, i.e., bypassing sandbox boundaries to access confidential data, or a poisoning attack, i.e., tricking a victim sandbox into transiently leaking its own memory content [34]. These attacks are particularly relevant to runtime environments that execute multiple (distrusting) tenants within a single thread. Securing runtime environments efficiently against breakout attacks is a major concern in industrial applications, e.g., serverless computing [40, 14].
@@
Although microarchitectural patches have found adoption in commercial CPUs to mitigate Meltdown and MDS attacks, Spectre attacks are still primarily addressed by countermeasures at the operating system and compiler levels. These mitigations entail a high performance overhead [42, 2] but are used to patch legacy systems nonetheless. This is primarily because secure speculation techniques in hardware, to date, necessitate significant modifications to the microarchitecture and memory subsystem (additionally incurring substantial performance overheads) which impede their adoption in industrial settings.
1-A Contribution
We propose Okapi – a novel hardware/software cross-layer architecture for TES attack mitigation that provides breakout and poisoning security with low performance and hardware overheads. At its core, the Okapi architecture extends the Page Table Entries (PTEs) in the Translation Look-aside Buffer (TLB) with an additional bit, the safe access bit, that marks a data page whenever one of its addresses is translated legally, i.e., non-speculatively, without causing an exception, in the currently running thread. If this bit is set, the CPU is allowed to load speculatively from addresses that belong to this page. On the other hand, Okapi blocks speculative execution of loads from pages that do not have the safe access bit set and thereby prevents formation of any transient execution side channel based on the content of these pages. Whenever there is a switch between software threads, all safe access bits in the TLB are reset to prevent the misuse of stale access rights. With these hardware features in place, Okapi removes all universal read gadgets and provides security against breakout attacks at thread level.
In addition, Okapi introduces a dedicated ISA instruction, called OkapiReset, to enable the programmer to remove the access rights explicitly, e.g., after sensitive data regions have been accessed legally. This allows a designer to prevent breakout attacks at a finer granularity, e.g., at function level.
Besides breakout attacks, Okapi also mitigates poisoning attacks [35]. Okapi introduces efficient hardware features to drastically limit the exploitable code space for Spectre gadgets. Whenever the instruction pointer speculatively crosses instruction page boundaries, Okapi prevents speculative read accesses, which inhibits the initial phase of a Spectre-BTB [28] attack. This measure prevents exploiting any gadgets outside the current trust domain and restricts the set of potential gadgets for a poisoning attack to only those that are located on the currently executing code page. With this additional hardware mitigation in place, the software measures elaborated in Sec. 4-B offer a trade-off between manual effort and performance to efficiently eliminate poisoning attacks in a given application.
@@ 1-A Contribution
  • We introduce Okapi, a new hardware/software cross-layer architecture to effectively and efficiently protect against breakout and poisoning attacks (cf. security objective in Sec. 3). In particular, Okapi eliminates all threats imposed by Spectre-like universal read gadgets with only minor hardware and performance overhead.
  • Okapi offers a set of incremental security features. The hardware features incur low performance and implementation overhead and provide breakout security for thread-sized sandboxes. They drastically limit the set of exploitable gadgets for poisoning attacks. Additional software measures allow for breakout security at a custom granularity below thread-level and full poisoning security.
  • Okapi introduces a total hardware overhead of less than 0.4 % additional logic and flipflops (cf. Sec. 8). For use cases without security requirements, the Okapi features can be switched off to avoid all performance overhead. The Okapi hardware features, when used without any software modification, incur only a small performance overhead of 3.17 % (cf. Sec. 7), while significantly improving security at a thread-level granularity. Hence, Okapi blocks all sandbox breakout attacks, e.g., through Spectre-PHT and Spectre-BTB.
  • Further security measures at the software level involving the OkapiReset instruction enable Okapi to deliver breakout security in custom-sized sandboxes, with competitive performance. We provide a case study for protecting mutually distrusting applications in a WebAssembly runtime environment, introducing only 1.68 % performance overhead.
  • Security against poisoning attacks can be achieved by identifying and hardening the gadgets only within the code base of the victim sandbox. Alternatively, confidential data can be accessed using a special load instruction, OkapiLoad, which does not mark the accessed page as safe for subsequent speculative loads and therefore prevents any speculative access to that page. This, although at the cost of additional performance overhead, relieves the software developer from manually identifying gadgets within the victim’s code base.
@@
The rest of the paper is organized as follows: Sec. 2 reviews background relevant for Okapi followed by Sec. 3 which formalizes the security objective. In Sec. 4 we introduce our Okapi concept and present its security implications in Sec. 5. Sec. 6 describes the Okapi architecture in detail before we evaluate different sandbox scenarios in a gem5 implementation of Okapi in Sec. 7. In Secs. 8 and 9, we analyze an RTL implementation of the Okapi hardware w.r.t. its hardware overhead and security guarantees. In addition, Sec. 9 inspects the size of trust domains in benchmark tests. Sec. 10 summarizes the state of the art of known mitigations for TES attacks.
2 Background
2-A Spectre Variants
@@
TES attacks [11] exploit features of modern processors like speculative or out-of-order execution to transiently access sensitive data in memory that would not have been accessed in the sequential execution of the program. Every attack involves three main phases:
  1. 1.
@@
    Setup: The adversary mistrains the targeted predictor and sets up the attacker process. Additional preparation like flushing the caches helps to enlarge the time window for the transient execution phase.
  2. 2.
@@
    Transient Access + Transmit: The victim code snippet is executed and transiently accesses the secret memory region. A subsequent transient operation transmits the information by altering the microarchitectural state of the processor in a secret-dependent way. After the core recognizes the misprediction, it rolls back to the last committed state, while some microarchitectural footprints (e.g., the cache state) remain in the system.
  3. 3.
@@
    Secret Recovery: The attacker recovers the secret data from the microarchitectural footprint and makes it visible in the architectural state, often using timing-based cache side channels like Flush+Reload [47] or a secret-dependent execution time in functional units [18].
The following part of this section briefly summarizes the three main classes of Spectre attacks:
@@
(a) Example pseudocode snippet of Spectre-PHT gadget
@@
(b) Example pseudocode snippet of Spectre-BTB gadget
@@
(c) Example pseudocode of Spectre-STL gadget
Figure 1: Spectre code gadgets (x86 pseudo-code)
Spectre-PHT
Also known as Spectre variant 1 [28], this attack exploits the pattern history table (PHT) that predicts the outcome of a branch decision. An example gadget is depicted in Fig. 1(a). During the setup phase, the adversary mistrains the predictor to predict the conditional ’jg’ jump to be not taken. Additionally, ARR_LEN is evicted from the cache to prolong the speculation time. Initiating the exploit, the attacker calls the code snippet with an out-of-bounds index rbx, targeting the memory address to be retrieved. Due to the mistraining, the processor assumes the provided index to be in bounds and executes both mov instructions. The gadget code transiently accesses the secret and stores it in the microarchitectural state of the data cache, by using the secret data as an address. After detecting the misprediction, the CPU reverts all program-visible effects of the transient instructions in the pipeline. However, the attacker can retrieve the data indirectly from the cache state. Because rbx is under the control of the attacker and can be of arbitrary value, the instruction sequence serves as a so-called universal read gadget.
Spectre-BTB/RSB
These variants, in contrast to Spectre-PHT, exploit speculation on indirect branches (variant 2 [28]) and return addresses (variant 5 [32, 29]). In this case, the predictors do not speculate on the outcome of a branch condition but on the target address the control flow jumps to. The attacker can poison the predictors to jump to any arbitrary gadget even within the victim’s code base to access the memory location of interest. Pseudocode for a Spectre v2 attack is depicted in Fig. 1(b). During the attack, the program counter makes a transient jump, leading to the execution of instructions that load secret data into the microarchitectural state before the misprediction is identified and corrected.
Spectre-STL
Besides control flow predictions that are used in Spectre variants 1 and 2, data flow predictions can open transient execution side channels as well ([1, 26]). Modern CPUs make predictions on target addresses within the load-store-unit to speed up data forwarding from one instruction to another. Not only does this pose a threat to data located in the store buffer, but it can also be used to construct a gadget that is able to transiently read from arbitrary memory addresses. An example of such a universal read gadget using data flow speculation is listed in Fig. 1(c). If the value of x is controlled by the attacker and the predictor is poisoned to predict the addresses addr0 and addr1 to match, the load-store-unit can forward the value of x to register rax. This value is then used as an address for the load instruction in line 3 to load from the attacker-controlled address x and afterwards transmit the content of this memory address through a microarchitectural side channel.
2-B Privilege Switches and Trust Domains
@@
A process is an instance of a possibly multithreaded, computer program that is being executed. Each process has an execution context, consisting of the threads’ program counters, register contents, address space and privilege level.
Threads
Threads are independent sequential execution paths in a program that run concurrently or quasi-parallel. We say quasi-parallel here because it depends on the number of available (logical) cores of the underlying CPU hardware whether two threads can be executed in parallel or whether the operating system has to schedule them on a single core. The current hardware state of each thread consists of an instruction pointer and the register contents.
Privilege Levels
The privilege level of a process defines its rights to modify other processes. Typically, the operating system and device drivers run with a higher privilege level than ordinary applications. Processes with a higher privilege level are allowed to modify processes with a lower privilege level but not vice versa. Such modifications include changes in the address space of a process or even changes in register and memory values. An address space is a set of virtual addresses that are readable, writable, or executable by the threads of the process [41].
@@
In this paper, a trust domain denotes the dynamically growing set of data pages that the currently running program has, so far, accessed legally and that is therefore allowed to be accessed speculatively. Depending on the chosen granularity of the sandboxes (e.g., threads, jobs, functions, etc.), trust domains can grow to different sizes. When handing over control from one sandbox to another all access permissions are reset. This ensures that sandboxes are prevented from accessing each other’s data pages.
2-C Virtual Memory and Paging
Most modern computer systems employ virtual memory (VM) [5], providing the user with an abstract programming interface. The key concept is to use virtual addresses that are translated by the VM system to actual physical addresses to access the requested data. Virtual memory is organized in pages [^1] One page is typically 4 KB in size. Modern VM systems usually also support 2 MB or 1 GB pages which are often called hugepages or megapages. that are mapped to physical page frames of the same size. The mapping for each process is stored in a separate page table that is managed by the operating system. If a process accesses a virtual address, the VM system performs a page table walk to translate the virtual address to a physical address. Page table walks are considered to be slow as they involve several accesses to physical memory to look up the mapping. Additionally, VM annotates pages with permissions (read, write, execute), thereby implementing simple memory protection and isolation.
The VM subsystem is crucial for system performance since about one third of fetched instructions interact with the memory system [5]. Our own experiments with the SPEC CPU2017 benchmark suite, in which up to 25 % of instructions are loads, support this statement (cf. Sec. 7-A). Therefore, computer architects introduce dedicated hardware to speed up address translation. The Translation Look-aside Buffer (TLB) is a low-latency cache that keeps track of recent address translations to avoid costly page table walks.
3 Security Objective
The security objective of Okapi is to prevent leakage of speculatively accessed data residing in data memory or in the cache hierarchy, i.e., data at rest. Similar to existing hardware mitigations against transient execution side channels, like STT [48], Eager/Naïve Delay [37], and NDA [45], Okapi targets the objective of weak speculative non-interference [24].
Following the terminology of [24], we label our memory locations either as high (H), i.e., confidential, or low (L), i.e., public.

### v3

@@
Abstract.
This paper introduces Okapi, a new hardware/software cross-layer architecture designed to mitigate Transient Execution Side Channel (TES) attacks, including Spectre variants, in modern computing systems. Okapi provides a hardware basis for secure speculation in sandboxed environments and can replace expensive speculation barriers in software.
@@ Abstract.
Okapi introduces the OkapiReset instruction for additional soft-ware-level security support. This instruction allows for fine-grained sandboxing with custom program sizes smaller than a thread, resulting in 2.34 % performance overhead in our WebAssembly runtime experiment.
@@
Spectre attacks, speculative execution, sandbox, system security, hardware design
journalyear: 2025copyright: acmlicensedconference: ACM Asia Conference on Computer and Communications Security; August 25–29, 2025; Hanoi, Vietnambooktitle: ACM Asia Conference on Computer and Communications Security (ASIA CCS ’25), August 25–29, 2025, Hanoi, Vietnamdoi: 10.1145/3708821.3733869isbn: 979-8-4007-1410-8/2025/08ccs: Security and privacy Systems securityccs: Security and privacy Side-channel analysis and countermeasuresccs: Hardware Hardware validation
1. Introduction
@@
Transient Execution Side Channel (TES) attacks, including numerous variants of Spectre (Kocher et al., 2019), persist as serious threats in modern computing systems. These vulnerabilities enable attackers to exploit side effects of speculative execution by tricking the processor into revealing information that is not accessed by the sequential semantics of the program. Since TES attacks exploit high-end features of modern microarchitectures such as speculative and out-of-order execution, eliminating them without introducing a significant performance penalty continues to be a major challenge. The most dangerous Spectre attacks involve so-called universal read gadgets that are capable of revealing the contents of any address in the entire memory space.
@@
A widely used software mechanism to restrict a program’s accessible resources is sandboxing (Wahbe et al., 1993). Untrusted code is often recommended to be executed within a sandbox (Google for Developers, 2018) in order to prevent it from stealing data or compromising the system (New York Times, 2019). As long as no additional measures are in place, Spectre-style attacks pose a severe threat to such software-level isolation techniques (Schwarzl et al., 2022). Attackers can either perform a breakout attack, i. e., bypassing sandbox boundaries to access confidential data, or a poisoning attack, i. e., tricking a victim sandbox into transiently leaking its own memory content (Narayan et al., 2021). These attacks are particularly relevant to runtime environments that execute multiple (distrusting) tenants within a single thread. Securing runtime environments efficiently against breakout attacks is a major concern in industrial applications, e. g., serverless computing (Schwarzl et al., 2022; Cloudflare, 2018).
@@
Although microarchitectural patches have found adoption in commercial CPUs to mitigate Meltdown and MDS attacks, Spectre attacks are still primarily addressed by countermeasures at the operating-system and compiler levels. These mitigations entail a high performance overhead (Turner, 2018; Amit et al., 2019) but are used to patch legacy systems nonetheless. This is primarily because secure speculation techniques in hardware, to date, necessitate significant modifications to the microarchitecture and the memory subsystem (additionally incurring substantial performance overheads) which impede their adoption in industrial settings.
1.1. Contribution
@@
  • We introduce Okapi, a new hardware/software cross-layer architecture that efficiently protects a system against all breakout attacks at a chosen granularity. With only 0.6 % additional cell area, Okapi achieves an exceptionally low performance overhead of only 2.34 % for securing a WebAssembly runtime.
  • Okapi also provides full protection against poisoning attacks, if certain assumptions for the software are fulfilled. In Sec. 4.2 we describe software measures to fulfill these assumptions and in Sec. 10 we discuss possible trade-offs between different levels of security and the corresponding overheads in hardware and performance.
  • In combination, the Okapi hardware and software measures are capable of mitigating TES attacks, ranging from a low-overhead countermeasure against Spectre-PHT up to providing weak speculative non-interference (Guarnieri et al., 2021).
At its core, the Okapi architecture extends the Page Table Entries in the Translation Look-aside Buffer (TLB) with an additional bit, the safe access bit, that marks a data page whenever one of its addresses is translated legally, i. e., non-speculatively, without causing an exception, in the currently running thread. If this bit is set, the CPU is allowed to speculatively load from addresses in this page. Conversely, Okapi blocks speculative execution of loads from pages that do not have the safe access bit set and thereby prevents formation of any transient execution side channel based on the content of these pages. Whenever there is a switch between software threads, all safe access bits in the TLB are reset to prevent the misuse of stale access rights. With these hardware features in place, Okapi removes all universal read gadgets (McIlroy et al., 2019) and provides security against breakout attacks at thread level.
In addition, Okapi introduces a dedicated ISA instruction, called OkapiReset, which allows the programmer to revoke access rights explicitly, for example, after sensitive data regions have been accessed legally. This allows a designer to prevent breakout attacks at a finer granularity, e. g., at function level.
Okapi’s protection mechanism guarantees that all data pages that cannot be read by any underlying software sandboxing mechanism will never be read speculatively by the microarchitecture either. This separation mechanism is orthogonal to existing in-process isolation techniques and efficiently closes their remaining gap for speculative execution.
In our case studies (Sec. 7), we demonstrate this for the state-of-the-art frameworks Wasmtime (Bytecode Alliance, 2017) and ERIM (Vahldiek-Oberwagner et al., 2019).
Besides breakout attacks, Okapi also mitigates poisoning attacks (Narayan et al., 2023). Okapi introduces efficient hardware features to drastically limit the exploitable code space for Spectre gadgets. Whenever the instruction pointer speculatively crosses instruction page boundaries, Okapi prevents speculative read accesses, which inhibits the initial phase of a Spectre-BTB (Kocher et al., 2019) attack. This measure prevents exploiting any gadgets outside the current trust domain and restricts the set of potential gadgets for a poisoning attack to only those that are located on the currently executing code page. With this additional hardware mitigation in place, the software measures elaborated in Sec. 4.2 offer a trade-off between manual effort and performance to efficiently eliminate poisoning attacks in a given application.
@@
The rest of the paper is organized as follows: Sec. 2 reviews background relevant for Okapi followed by Sec. 3 which formalizes the security objective. In Sec. 4 we introduce our Okapi concept and present its security implications in Sec. 5. Sec. 6 describes the Okapi architecture in detail before we evaluate different sandbox scenarios in a gem5 implementation of Okapi in Sec. 7. In Secs. 8 and 9, we analyze an RTL implementation of the Okapi hardware w. r. t. its hardware overhead and security guarantees. In addition, Sec. 9 inspects the size of trust domains in benchmark tests. Sec. 10 presents “sweet spots” for a practical application of Okapi and Sec. 11 summarizes the state of the art of known mitigations for TES attacks.
2. Background
2.1. Spectre Variants
@@
A TES attack (Canella et al., 2019) involves three main phases to transiently access sensitive data in memory that would not have been accessed in the sequential execution of the program:
Setup:
@@
The adversary mistrains the targeted predictor and sets up the attacker process. Additional preparation like flushing the caches helps to enlarge the time window for the transient execution phase.
Transient Access + Transmit:
@@
The victim code snippet is executed and transiently accesses the secret memory region. A subsequent transient operation transmits the information by altering the microarchitectural state of the processor in a secret-dependent way. After the core recognizes the misprediction, it rolls back to the last committed state, while some microarchitectural footprints (e. g., the cache state) remain in the system.
Secret Recovery:
@@
The attacker recovers the secret data from the microarchitectural footprint and makes it visible in the architectural state, often using timing-based cache side channels like Flush+Reload (Yarom and Falkner, 2014) or a secret-dependent execution time in functional units (Fadiheh et al., 2020).
Different types of Spectre attacks utilize different sources of speculative execution and therefore mainly vary in the second phase.
Spectre-PHT (Kocher et al., 2019), also known as Spectre variant 1, exploits the pattern history table (PHT) that predicts the outcome of a branch. An example gadget is depicted in Fig. 1(a). If the offset of the transient read instruction (rbx in Fig. 1(a)) is controlled by the attacker, the instruction sequence serves as a so-called universal read gadget that allows the attacker to read from the entire address space.
Spectre-BTB and Spectre-RSB exploit speculation on indirect branches (variant 2 (Kocher et al., 2019)) and return addresses (variant 5 (Maisuradze and Rossow, 2018; Koruyeh et al., 2018)). In this case, the predictors do not speculate on the outcome of the decision of a branch but on its target address. The attacker can poison the predictors to jump to an arbitrary gadget even within the victim’s code base to access the memory location of interest. Pseudocode for a Spectre v2 attack is depicted in Fig. 1(b).
Besides control flow predictions that are used in Spectre variants 1 and 2, data flow predictions can open transient execution side channels as well (AMD, 2019; Horn, 2018). Modern CPUs make predictions on target addresses within the load-store-unit to speed up data forwarding from one instruction to another. Not only does this pose a threat to data located in the store buffer, but it can also be used to construct a gadget that is able to transiently read from arbitrary memory addresses. An example of such a universal read gadget using data flow speculation is listed in Fig. 1(c). If the value of x is controlled by the attacker and the predictor is poisoned to predict the addresses addr0 and addr1 to match, the load-store-unit can forward the value of x to register rax. This value is then used as an address for the load instruction in line 3 to load from the attacker-controlled address x and afterward transmit the content of this memory address through a microarchitectural side channel.
@@
[…]
Example pseudocode snippet of a Spectre-PHT gadget. An array access is guarded by an if-statement that ensures the array index to be in-bounds. If the outcome of the if-condition is predicted to be in-bounds, then the array access is performed and the returned value get’s encoded into the cache to leak it.
(a) Example pseudocode snippet of a Spectre-PHT gadget
@@
[…]
Example pseudocode snippet of a Spectre-BTB gadget. An attacker controlls the target prediction for indirecct jumps and can use this to speculatively divert the controll flow to a cache encoder gadget to leak the secret from a register.
(b) Example pseudocode snippet of a Spectre-BTB gadget
@@
[…]
Example pseudocode of a Spectre-STL gadget. The attacker controlls an input value that is stored at some address addr0. If the CPU predicts addr0 to be equal to some addr1 that the code loads from, it speculatively forwards the attacker controlled value to the load instruction. If this loaded value is used as an index, the attacker can make the code speculatively read a secret value that can be leaked through, e. g., the cache.
(c) Example pseudocode of a Spectre-STL gadget
Figure 1. Spectre code gadgets (x86 pseudo-code)
[…]
Spectre code gadgets are are displayed in x86 pseudo-code. Subfigure a shows a Spectre-PHT gadget. Subfigure b shows a Spectre-BTB gadget. Subfigure c shows a Spectre-STL gadget.
2.2. Privilege Switches and Trust Domains
@@
A process is an instance of a possibly multithreaded computer program that is being executed. Each process has an execution context, consisting of an address space, a privilege level, and a list of threads. Threads are independent sequential execution paths in a process that run concurrently if sufficiently many (logical) CPU cores are available. The current hardware state of each thread consists of an instruction pointer and the register contents. An address space is a set of virtual addresses that are readable, writable, or executable by the threads of the process (Tanenbaum and Bos, 2014). The privilege level of a process defines its rights to modify other processes. Processes with a higher privilege level are allowed to modify processes with a lower privilege level but not vice versa. Such modifications include changes in the address space of a process or even changes in register and memory values.
@@
In this paper, a trust domain denotes the currently executed code page together with the dynamically growing set of data pages that the currently running thread has, so far, accessed legally and that is therefore allowed to be accessed speculatively. Depending on the chosen granularity of the sandboxes (e. g., threads, jobs, functions, etc.), trust domains can grow to different sizes. When handing over control from one sandbox to another all access permissions are reset. This ensures that sandboxes are prevented from accessing each other’s data pages.
2.3. Virtual Memory and Paging
Most modern computer systems employ virtual memory (VM) (Bhattacharjee and Lustig, 2017), providing the user with an abstract programming interface. The key concept is to use virtual addresses that are translated by the VM system to physical addresses to access the requested data. Virtual memory is organized in pages [^1] One page is typically 4 KB in size. Modern VM systems usually also support 2 MB or 1 GB pages which are often called hugepages or megapages. that are mapped to physical page frames of the same size. The mapping for each process is stored in a separate page table that is managed by the operating system. If a process accesses a virtual address, the VM system performs a page table walk to translate the virtual address to a physical address. Page table walks are considered to be slow as they involve several accesses to physical memory to look up the mapping. Additionally, VM annotates pages with permissions (read, write, execute), thereby implementing simple memory protection and isolation. The VM subsystem is crucial for system performance since about one third of fetched instructions interact with the memory system (Bhattacharjee and Lustig, 2017). Our own experiments with the SPEC CPU2017 benchmark suite, in which up to 25 % of instructions are loads, support this statement (cf. Sec. 7.1). Therefore, computer architects introduce dedicated hardware to speed up address translation. The Translation Look-aside Buffer (TLB) is a low-latency cache that keeps track of recent address translations to avoid costly page table walks.
3. Security Objective
The security objective of Okapi is to prevent leakage of data residing in data memory or in the cache hierarchy, i. e., data at rest. Similar to existing hardware mitigations against transient execution side channels, like STT (Yu et al., 2019), Eager/Naïve Delay (Sakalis et al., 2019), and NDA (Weisse et al., 2019), Okapi targets the objective of weak speculative non-interference (Guarnieri et al., 2021).
Following the terminology of (Guarnieri et al., 2021), we denote a memory location as high (H) if it stores confidential data, or as low (L) if the data at the location is public.